# Effects of Collaboration on the Performance of Interactive Theme Discovery Systems

### v2

@@ 8 Limitations
(2) As prior work has shown Hoyle et al. (2021), automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
@@ Appendix A Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as Fang et al. (2021), which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
@@ B.1 Relational Approach Operational Details
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin and Han (2010) and Hierarchical Density-Based Clustering Mendelsohn et al. (2021) algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots van der Maaten and Hinton (2008).
@@
Table 3: Interactive Operations for the Pacheco et al. (2023) System
@@ Appendix D LLM-based Experimental Configuration and Prompt Details
code-book Gwet’s $AC_{1}$ Unlabeled Docs
@@
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (Grattafiori et al., 2024) model for all generation tasks:

### v3

@@ 8 Limitations
(2) As prior work has shown 10.5555/3540261.3540416, automated metrics and human evaluation do not always correlate well for unsupervised or semi-supervised theme induction tasks like topic modeling. Although we did incorporate two types of human evaluation: a post-hoc manual verification (Sec 5) and a user study (Sec 6), both of these evaluations were limited in scope. This is primarily due to how expensive and time consuming it is to perform an exhaustive human evaluation. While we are confident in the findings of our study, a larger study could further strengthen our conclusions.
@@ Appendix A Topic Model Experimental Settings
To ensure comparable results, we use the same starting distribution of 13 topics for all our experiments using the same hyperparameters as fang-etal-2021-query, which are $\alpha=1.0$, $\beta=0.5$, $\gamma=1.5$. The QDTM also allows queries to be input prior to topic model initialization to produce partitions that follow prior knowledge, but we do not take advantage of this capability. The same initial topic model is provided for all experiments.
@@ B.1 Relational Approach Operational Details
Finding Partitions Experts can find partitions in the space of unassigned instances. We currently support the K-means Jin2010 and Hierarchical Density-Based Clustering mendelsohn-etal-2021-modeling algorithms.
@@
Visualizing Global Explanations Experts can visualize aggregated statistics and explanations for the global state of the system. To do this, we aggregate all instances in the database. Explanations include theme distribution, coverage statistics, and t-sne plots JMLR:v9:vandermaaten08a.
@@
Table 3: Interactive Operations for the pacheco-2023-interactive System
@@ Appendix D LLM-based Experimental Configuration and Prompt Details
code-book Gwet’s $AC_{1}$ # Unlabeled Docs
@@
Using the LLM, we generate 3 additional templates based on a prompt structure provided in the original work. We used the Llama 3.2 3B-Instruct (grattafiori2024llama3herdmodels) model for all generation tasks: